OpenStudy (anonymous):

What is the constant sum of the ellipse with foci at (-12, 0) and (12, 0) that passes through the point (-15, 0)?

OpenStudy (dumbcow):

constant sum ? you mean equation of the ellipse...

OpenStudy (anonymous):

constant sum. That's what the equation gives me

OpenStudy (dumbcow):

 looks like major vertices are at (-15,0) , (15,0) a = 15 c = 12 a^2 = b^2 + c^2 225 = b^2 + 144 b = 9 equation: \[\frac{x^{2}}{225}+\frac{y^{2}}{81} = 1\]